OpenStudy (anonymous):

verify the identity. show all work. tanθ cosθ-sin^2 θ=cos^2 θ

OpenStudy (abb0t):

\(tan(\theta) = \frac{sin(\theta}{cos(\theta)}\)

OpenStudy (anonymous):

what?

OpenStudy (abb0t):

that's one of the identities you need to know, also \(sin^2(\theta) = 1-cos^2(\theta)\)

OpenStudy (dan815):